Solve the initial value problem "2f''(t)-3f'(t)+f(t)=te^t, f(0)=1,f'(0)=\frac{3}{2}"


Using the point transformation [y = exp(ψ)], transform the equation (Eହ


) into a first

order non-linear O.D.E with independent variable x and dependent variable w, whereby w ≔ ψ

.


Afterwards, use the second point transformation ቂw =

ିଶఉᇲ(௫)

ቃ to change this first order O.D.E to a

second order, linear and homogeneous O.D.E with independent variable x and dependent variable β .


1
Expert's answer
2020-05-25T20:50:17-0400

The solution consists of two parts:



"f(t)=f_{hom}(t)+f_{par}(t)\\[0.3cm]
f_{hom}(t)-\text{solution of a homogeneous equation}\\[0.3cm]
2f''(t)-3f'(t)+f(t)=0\\
f_{par}(t)-\text{particular solution of an inhomogeneous equation}\\[0.3cm]
2f''(t)-3f'(t)+f(t)=t\cdot e^t"

1 STEP: We solve the homogeneous equation



"2f''(t)-3f'(t)+f(t)=0"

We will seek a solution in the form "f(t)=e^{kt}" , then



"f(t)=e^{kt}\longrightarrow
\left\{\begin{array}{l}
f'(t)=k\cdot e^{kt}\\[0.3cm]
f''(t)=k^2\cdot e^{kt}
\end{array}\right.\longrightarrow\\[0.3cm]
2\cdot k^2\cdot e^{kt}-3\cdot k\cdot e^{kt}+e^{kt}=0\\[0.3cm]
e^{kt}\cdot\left(2k^2-3k+1\right)=0\longrightarrow\underbrace{2k^2-3k+1=0}_{\text{quadratic equation}}\\[0.3cm]
\sqrt{D}=\sqrt{\left(-3\right)^2-4\cdot 2\cdot1}=\sqrt{1}=1\\[0.3cm]
\left[\begin{array}{l}
k_1=\displaystyle\frac{-(-3)+1}{2\cdot2}=1\\[0.3cm]
k_2=\displaystyle\frac{-(-3)-1}{2\cdot2}=\displaystyle\frac{1}{2}
\end{array}\right.\\[0.3cm]
\boxed{f_{hom}(t)=C_1\cdot e^t+C_2\cdot e^{t/2}}"

2 STEP: We solve the inhomogeneous equation



"2f''(t)-3f'(t)+f(t)=t\cdot e^t"

Minimum required theory: when solving inhomogeneous equations of the form



"F\left(f,f',f'',f''',\ldots\right)=g(t)"

if the right side is "g(t)=P_m(t)\cdot e^{\alpha t},\,\,\text{where}\,\,\, P_m(t)-\text{polynomial}" , then to look for the particular solution in the form



"\left[\begin{array}{l}
\widetilde{P_m}(t)\cdot e^{\alpha t}-\alpha\,\,\text{not the root of the characteristic equation}\\[0.3cm]
x^s\widetilde{P_m}(t)\cdot e^{\alpha t}-\alpha\,\,\text{the root of the characteristic equation}
\end{array}\right."

In our case,



"g(t)=t\cdot e^{1\cdot t}\longrightarrow k_1=1-\text{multiplicity root 1}\\[0.3cm]
f_{par}(t)=t\cdot\left(At+B\right)\cdot e^t\equiv\left(At^2+Bt\right)\cdot e^t\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
f'(t)=\left(At^2+Bt\right)\cdot e^t+\left(2At+B\right)\cdot e^t\\[0.3cm]
f'(t)=\left(At^2+(2A+B)t+B\right)\cdot e^t\\[0.3cm]
f''(t)=\left(At^2+(2A+B)t+B\right)\cdot e^t+\left(2At+(2A+B)\right)\cdot e^t\\[0.3cm]
f''(t)=\left(At^2+(4A+B)t+(2A+2B)\right)\cdot e^t
\end{array}\right."

Substitute in the original equation



"2\left(At^2+(4A+B)t+(2A+2B)\right)\cdot e^t-\\[0.3cm]
-3\left(At^2+(2A+B)t+B\right)\cdot e^t+\left(At^2+Bt\right)\cdot e^t=t\cdot e^t\\[0.3cm]
(2A-3A+A)t^2+(2(4A+B)-3(2A+B)+B)t+\\[0.3cm]
+(2(2A+2B)-3B)=t\\[0.3cm]
0\cdot t^2+(2A)\cdot t+(4A+B)=t\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
2A=1\\[0.3cm]
4A+B=0
\end{array}\right.
\longrightarrow
\left\{\begin{array}{l}
A=\displaystyle\frac{1}{2}\\[0.3cm]
B=-4A=-2
\end{array}\right."

Conclusion,



"\boxed{f_{par}(t)=\left(\frac{t^2}{2}-2t\right)\cdot e^t}"

General solution,



"\boxed{f(t)=C_1\cdot e^t+C_2\cdot e^{t/2}+\left(\frac{t^2}{2}-2t\right)\cdot e^t}"

The constants are determined from the initial conditions



"f(0)=1=C_1\cdot e^0+C_2\cdot e^{0/2}+\left(\frac{0^2}{2}-2\cdot0\right)\cdot e^0\\[0.3cm]
\boxed{C_1+C_2=1}\\[0.3cm]
f'(t)=C_1\cdot e^t+\frac{C_2}{2}\cdot e^{t/2}+(2t-2)\cdot e^t+\left(\frac{t^2}{2}-2t\right)\cdot e^t\\[0.3cm]
f'(0)=\frac{3}{2}=C_1\cdot e^0+\frac{C_2}{2}\cdot e^{0/2}+(2\cdot0-2)\cdot e^0+\left(\frac{0^2}{2}-2\cdot0\right)\cdot e^0\\[0.3cm]
\boxed{C_1+\frac{C_2}{2}-2=\frac{3}{2}}"

It remains to solve the system of equations



"\left\{\begin{array}{l}
C_1+C_2=1\\[0.3cm]
\left.C_1+\displaystyle\frac{C_2}{2}-2=\displaystyle\frac{3}{2}\right|\times(2)
\end{array}\right.\longrightarrow\left\{\begin{array}{l}
C_2=1-C_1\\[0.3cm]
2C_1+C_2-4=3\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
C_2=1-C_1\\[0.3cm]
2C_1+1-C_1-4=3\end{array}\right.\longrightarrow
\left\{\begin{array}{l}
C_2=-5\\[0.3cm]
C_1=6\end{array}\right."


Conclusion, solution of IVT is



"\boxed{f(t)=6\cdot e^t-5\cdot e^{t/2}+\left(\frac{t^2}{2}-2t\right)\cdot e^t}"

ANSWER



"f(t)=6\cdot e^t-5\cdot e^{t/2}+\left(\frac{t^2}{2}-2t\right)\cdot e^t"